OpenStudy (anonymous):

Hey does anyone need help in math? I have successfully completed Alg 1 and 2 and geometry with either all A's or A's and B's if you need help with any classes I can explain problems to you and such.

OpenStudy (anonymous):

Screenshot your question and then click this link and click on the speech bubble and talk to me there. I will walk you through the problem(s) if I remember how to do them

OpenStudy (anonymous):

me

OpenStudy (mony01):

Do you know how to do this problem? Consider the curve x^2+4xy+y^2=13. Which one of these three points is on this curve: (1,1),(2,1),(1,-2) and why? Find slope of the tangent line to the curve at that point.

OpenStudy (anonymous):

Oh my now way. That sounds like the stuff I was working on in precalc. I'm going to a tutor to get help with stuff like this tomorrow actually Sorry.

OpenStudy (anonymous):

@bmeyer74 what do you need help with?

OpenStudy (anonymous):

@mony01 Just substitute the points into the equation and see if it works. Remember a point is (x, y), so for example: (1, 1) 1^2 + 4(1)(1) + (1)^2 = 13 1 + 4 + 1 = 13 6 = 13 This is not true Therefore, (1,1) is not on the curve. I'm sure you can do the rest!

OpenStudy (aravindg):

Very innovative :)

OpenStudy (mony01):

@Pompeii00 thanks but when it says to find the slope of tangent line, what do I do there?

OpenStudy (anonymous):

@mony01 that starts to be calculus...

hartnn (hartnn):

@mony01 you need to differentiate x^2+4xy+y^2=13. first, implicitly, do you know what that is ?